# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 6.3. The finiteness property
Let $\pi\colon\mathcal{X}\to\Delta$ be a quasi-projective open Kulikov degeneration. Then, there exists a quasi-projective open Kulikov degeneration $\pi^{\prime}\colon\mathcal{X}^{\prime}\rightarrow\Delta$, with affinization $\overline{\mathcal{X}}$ and having a good effective divisor.
@@
The result follows immediately from Lemma 6.12, and from the fact that, if $h:\mathcal{X}\rightarrow\mathcal{Z}$ is an isomorphism in codimension one, then the exceptional locus of $g$ is a configuration of curves as in Definition 6.8 i). ∎
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}$, there exists finitely many $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$.
@@
By Lemma 6.13, there exists a quasi-projective open Kulikov degeneration $\mathcal{X}^{\prime}\rightarrow\Delta$ with affinization $\overline{\mathcal{X}}$, and with a good effective divisor as in Definition 6.8. In particular, it follows by Theorem 6.11 that Lemma 6.14 holds for $\mathcal{X}^{\prime}\rightarrow\Delta$. As $\mathcal{X}^{\prime}$ and $\mathcal{X}$ are related by a sequence of M1 and M2 flops, Lemma 6.14 holds for $\mathcal{X}\rightarrow\Delta$ by Theorems 6.2-6.4. ∎
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. For every $p,q,r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$, we have $h(r)=h(p)+h(q)$.
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Then, for every $p,q\in C(P)_{\mathbb{Z}}$, there exists finitely many $r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$.
@@
6.4. Polarized mirrors and the movable secondary fan
6.4.1. The polarized mirror family to an open Kulikov degeneration
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. It follows from Theorem 6.16, that the formula (5.1) defines an algebra structure on the $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$-module
$\mathcal{R}_{\mathcal{X}}=\bigoplus_{p\in C(P)_{\mathbb{Z}}}\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]\,\vartheta_{p}\,.$
@@ 6.4.1. The polarized mirror family to an open Kulikov degeneration
$\mathcal{Y}_{\mathcal{X}}:=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}\longrightarrow S_{\mathcal{X}}:=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$
endowed with the line bundle $\mathcal{L}_{\mathcal{X}}:=\mathcal{O}_{\mathcal{Y}_{\mathcal{X}}}(1)$.
For a quasi-projective open Kulikov degeneration $\mathcal{X}\rightarrow\Delta$, the contraction $\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is a relative Mori dream space by Proposition 3.36. The corresponding Mori fan $MF(\mathcal{X}/\overline{\mathcal{X}})$ is a complete fan in $\mathrm{Pic}(\mathcal{X}/\overline{\mathcal{X}})_{\mathbb{R}}$, whose maximal dimensional cones are indexed by the finitely many $\mathbb{Q}$-factorial birational models $f^{\prime}:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ of $\mathcal{X}$ over $\overline{\mathcal{X}}$, and are given by $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})+\langle\mathrm{Ex}(f^{\prime})\rangle$, where $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ is the relative nef cone of $\mathcal{X}^{\prime}$ over $\overline{\mathcal{X}}$, and $\langle\mathrm{Ex}(f^{\prime})\rangle$ is the non-negative span of the $f^{\prime}$-exceptional divisors.
If $\langle\mathrm{Ex}(f^{\prime})\rangle\neq\{0\}$, that is, if at least one divisor is contracted, the corresponding cone in $MF(\mathcal{X}/\overline{\mathcal{X}})$ is referred to as bogus. In other words, a cone is non-bogus if $\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ is small, that is, by Lemma 6.1, if $\mathcal{X}^{\prime}\rightarrow\Delta$ is a projective resolution, and $\mathcal{X}^{\prime}\rightarrow\Delta$ is an open Kulikov degeneration. The union of non-bogus cones is the movable cone $\mathrm{Mov}(\mathcal{X}/\overline{\mathcal{X}})$.
Following [50], we define the secondary fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$ as a coarsening of the Mori fan $MF(\mathcal{X}/\overline{\mathcal{X}})$. We first define the movable secondary fan $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$ as a coarsening of the restriction $\mathrm{MovF}(\mathcal{X}/\overline{\mathcal{X}})$ of the Mori fan to the movable cone $\mathrm{Mov}(\mathcal{X}/\overline{\mathcal{X}})$. Two maximal cones $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ of $\mathrm{MovF}(\mathcal{X}/\overline{\mathcal{X}})$ are called equivalent if the corresponding multivalued PL functions $\varphi^{\prime}$ and $\varphi^{\prime\prime}$ as in §6.2 are equal.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. There exists a unique fan $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$ with support the movable cone $\mathrm{Mov}(\mathcal{X}/\overline{\mathcal{X}})$, whose maximal cones are the unions of equivalent maximal cones of $\mathrm{MovF}(\mathcal{X}/\overline{\mathcal{X}})$.
@@
The proof is analogous to the proof of Theorem [50, Theorem 2.12]. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Two maximal cones $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ of $\mathrm{Mov}(\mathcal{X}/\overline{\mathcal{X}})$ are equivalent if and only if $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are connected by a sequence of M1 flops.
@@
If $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are connected by a sequence of M1 flops, then $\mathcal{X}_{0}^{\prime}$ and $\mathcal{X}_{0}^{\prime\prime}$ have the same double curves, so we have $\varphi^{\prime}=\varphi^{\prime\prime}$, and the maximal cones $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ are equivalent.
Conversely, if $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ are equivalent, then, by Lemma 6.18, $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ are both contained in a maximal cone $C$ of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$. As $C$ is convex, there exists a path contained in $C$, starting in $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$, ending $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$, and only intersecting cones of $MF(\mathcal{X}/\overline{\mathcal{X}})$ of codimension zero or one. Therefore, it is enough to prove the result when $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime\prime}/\overline{\mathcal{X}})$ share a common codimension one face, that is, when $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are related by an elementary flop, which by Lemma 6.1 is either a M1 or a M2 flop. By Lemma 6.5, $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ would not be equivalent if they were related by an M2 flop, and so we conclude that $\mathcal{X}^{\prime}$ and $\mathcal{X}^{\prime\prime}$ are in fact related by an M1 flop. ∎
Let $C$ be a maximal cone of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$. By Theorem 6.6, all mirror algebras $\mathcal{R}_{\mathcal{X}^{\prime}}$ with $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})\subset C$ have the same structure constants $N_{pqr}^{\beta}(\mathcal{X}^{\prime})$. In particular, these structure constants are only non-zero if $\beta\in NE(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ for every $\mathcal{X}^{\prime}$ with $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})\subset C$, that is if $\beta\in C^{\vee}_{\mathbb{Z}}$, where $C^{\vee}$ is the dual cone of $C$, and so define a $\mathbb{C}[C^{\vee}_{\mathbb{Z}}]$-algebra structure on
We denote
$\mathcal{Y}_{C}:=\mathrm{Spec}\,\mathcal{R}_{C}\longrightarrow\mathrm{Spec}\,\mathbb{C}[C^{\vee}_{\mathbb{Z}}]\,$
and $\mathcal{L}_{C}:=\mathcal{O}_{\mathcal{Y}_{C}}(1)$.
If $C$ and $C^{\prime}$ are two maximal cones of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$, with corresponding multivalued PL functions $\varphi_{C}$ and $\varphi_{C^{\prime}}$, then it follows from Theorem 6.6 that the map
$\vartheta_{p}$ $\longmapsto t^{\varphi_{C}(p)-\varphi_{C^{\prime}}(p)}\vartheta_{p}\,,$
is an isomorphism of $\mathbb{C}[(C\cap C^{\prime})^{\vee}_{\mathbb{Z}}]$-algebras. If $C$, $C^{\prime}$, and $C^{\prime\prime}$ are three maximal cones of $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$, then the cocycle condition $\psi_{C,C^{\prime\prime}}=\psi_{C,C^{\prime}}\circ\psi_{C^{\prime},C^{\prime\prime}}$ holds. Therefore, using the isomorphisms $\psi_{C,C^{\prime}}$, one can glue the families $(\mathcal{Y}_{C},\mathcal{L}_{C})\rightarrow\mathrm{Spec}\,\mathbb{C}[C^{\vee}_{\mathbb{Z}}]$ into a polarized family
$\nu_{\mathrm{msec}}:(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}})\longrightarrow S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ (6.5)
over the toric variety $S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ defined by the movable secondary fan $\mathrm{MovSec}(\mathcal{X}/\overline{\mathcal{X}})$. By Equation (6.4), for every $p\in(hP)_{\mathbb{Z}}$, $\vartheta_{p}$ naturally becomes a section of the line bundle $(\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}})^{\otimes h}\otimes\nu_{\mathrm{msec}}^{\star}L_{p}$ on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$, where $L_{p}$ is the line bundle on $S^{\mathrm{\mathrm{msec}}}_{\mathcal{X}}$ defined by the 1-cocycle $t^{\varphi_{C}(p)-\varphi_{C^{\prime}}(p)}$.
Our goal in the remaining part of this section is to describe divisors $\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}$ and $\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}$, and show that the pair $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}},\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}})$ forms a KSBA stable family. We first describe $\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}$ below.
@@
For every $p\in P_{\mathbb{Z}}$, the line bundle $L_{p}$ is trivial, that is, $\vartheta_{p}$ is naturally a section of $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$.

### v3

@@ 6.3. The finiteness property
Let $\pi\colon\mathcal{X}\to\Delta$ be a quasi-projective open Kulikov degeneration. Then, there exists a quasi-projective open Kulikov degeneration $\pi^{\prime}\colon\mathcal{X}^{\prime}\rightarrow\Delta$, with affinization $\mathcal{X}^{\mathrm{can}}$, and admitting a good effective divisor.
@@
The result follows immediately from Lemma 6.20, and from the fact that, if $g:\mathcal{X}\rightarrow\mathcal{Z}$ is an isomorphism in codimension one, then the exceptional locus of $g$ is a configuration of curves as in Definition 6.16 i). ∎
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}$, there exists finitely many $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$.
@@
By Lemma 6.21, there exists a quasi-projective open Kulikov degeneration $\mathcal{X}^{\prime}\rightarrow\Delta$ with affinization $\mathcal{X}^{\mathrm{can}}$, and with a good effective divisor as in Definition 6.16. In particular, it follows by Theorem 6.19 that Lemma 6.22 holds for $\mathcal{X}^{\prime}\rightarrow\Delta$. Hence, Lemma 6.22 also holds for $\mathcal{X}\rightarrow\Delta$ by Theorem 6.14. ∎
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. For every $p,q,r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$, we have $h(r)=h(p)+h(q)$.
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Then, for every $p,q\in C(P)_{\mathbb{Z}}$, there exists finitely many $r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$.
@@
6.4. Polarized mirror family to an open Kulikov degeneration
6.4.1. Polarized mirror family
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration with tropicalization $(C(P),C(\mathscr{P}))$. It follows from Theorem 6.24 that the product of theta functions given by Equation (5.1) defines an algebra structure on the $\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$-module
$\mathcal{R}_{\mathcal{X}}=\bigoplus_{p\in C(P)_{\mathbb{Z}}}\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]\,\vartheta_{p}\,.$
@@ 6.4.1. Polarized mirror family
$\mathcal{Y}_{\mathcal{X}}:=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}\longrightarrow S_{\mathcal{X}}:=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$
endowed with the sheaf $\mathcal{L}_{\mathcal{X}}:=\mathcal{O}_{\mathcal{Y}_{\mathcal{X}}}(1)$.
6.4.2. Torus action
To prove properties of the polarized mirror family $(\mathcal{Y}_{\mathcal{X}},\mathcal{L}_{\mathcal{X}})$, we will make use as in [44, 53, 70] of a torus action, defined as follows. Denote by $T_{0}$ the torus $\mathrm{Spec}\,\mathbb{C}[\mathbb{Z}^{P_{\mathbb{Z}}}]$. The character lattice $\mathbb{Z}^{P_{\mathbb{Z}}}$ of $T_{0}$ can be naturally identified with the group of divisors of $\mathcal{X}$ supported on $\mathcal{X}_{0}$, since the points of $P_{\mathbb{Z}}$ are in a natural one-to-one correspondence with the irreducible components of $\mathcal{X}_{0}$. In particular, for every divisor $F$ supported on $\mathcal{X}_{0}$, we have a corresponding one-parameter subgroup $\mathbb{C}^{\star}\subset T_{0}$. There is a natural action of $T_{0}$ on $S_{\mathcal{X}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$ such that, for every divisor $F$ supported on $\mathcal{X}_{0}$, the corresponding one-parameter subgroup acts with weights $F\cdot\beta$ on the monomials $t^{\beta}$.
@@ 6.4.2. Torus action
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. There exists a unique action of the torus $T_{0}=\mathrm{Spec}\,\mathbb{C}[\mathbb{Z}^{P_{\mathbb{Z}}}]$ on the mirror family
$\mathcal{Y}_{\mathcal{X}}=\mathrm{Proj}\,\mathcal{R}_{\mathcal{X}}\longrightarrow S_{\mathcal{X}}=\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]\,$,
@@
This follows immediately from Lemma 6.15– see also [51, Construction 1.24] in the general context of the intrinsic mirror construction. ∎
6.4.3. Central fiber, finite generation, and theta divisor
We first describe the central fiber of the polarized mirror family. Then, using the action of the torus $T_{0}$, we prove the finite generation of the mirror algebra, and we show that the sheaf $\mathcal{L}_{\mathcal{X}}$ is actually a relatively ample line bundle on the mirror family $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. Finally, we define a natural theta divisor $\mathcal{C}_{\mathcal{X}}\in|\mathcal{L}_{\mathcal{X}}|$.
@@ 6.4.3. Central fiber, finite generation, and theta divisor
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration with tropicalization $(C(P),C(\mathscr{P}))$. Let $(Y_{0_{\mathcal{X}}},L_{0,\mathcal{X}})$ be the central fiber of the polarized mirror $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ over the unique torus fixed point $0_{\mathcal{X}}$ of the affine toric variety $S_{\mathcal{X}}$. Then, the connected components $(Y_{0_{\mathcal{X}}}^{v},L_{0,\mathcal{X}}^{v})$ of the normalization of $(Y_{0_{\mathcal{X}}},L_{0,\mathcal{X}})$ are in one-to-one correspondence with the integral points $v\in P_{\mathbb{Z}}$ and are all isomorphic to $(\mathbb{P}^{2},\mathcal{O}_{\mathbb{P}^{2}}(1))$. Moreover, $(Y_{0_{\mathcal{X}}}^{v},L_{0,\mathcal{X}}^{v})$ are glued to form $(Y_{0_{\mathcal{X}}},L_{0,\mathcal{X}})$ in such a way that the resulting intersection complex is given by $(P,\mathscr{P})$.
@@
This follows from [48, §2.1], where the central fibers of mirror families given by the intrinsic mirror construction are described explicitly. ∎
@@
Let $\pi:\mathcal{X}\to\Delta$ be a generic quasi-projective open Kulikov degeneration and let $F$ be a good effective divisor on $\mathcal{X}$. Let $G$ be the face of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ spanned by the finitely many curves $C$ such that $F\cdot C=0$, and $J=NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\setminus G$. Then, the restriction of the family $\mathcal{Y}_{\mathcal{X}}\to\mathcal{S}_{\mathcal{X}}$ to the toric stratum
$\mathrm{Spec}\,\mathbb{C}[G]\cong\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]/J\subset S_{\mathcal{X}}$
is a Mumford toric degeneration with central fiber $Y_{0_{\mathcal{X}}}$.